OpenStudy (anonymous):

How do you prove this by induction...? Like step two after you sub in k+1 ? (1+x)^2 is greater than or equal to 1+nx where x>-1

OpenStudy (anonymous):

must be a typo here is there some other \(n\)?

OpenStudy (anonymous):

whoops sorry the ^2 is a ^n

OpenStudy (anonymous):

for example, if \(n=100\) and \(x=2\) then there is no way that \[(1+2)^2\geq 1+100\times2\]

OpenStudy (anonymous):

(1+x) ^n greater or equal to 1+ nk where x>-1

OpenStudy (anonymous):

kk that makes more sense

OpenStudy (anonymous):

you know how to do a proof by induction? this should not be hard if you know it

OpenStudy (anonymous):

first you have to show it is true if \(n=1\) i.e. you have to show that \[(1+x)^1\geq 1+1\times x\] which is pretty darn obvious because they are the same ok so far?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

uh i got up until k+1

OpenStudy (anonymous):

ok hold on first you get to ASSUME it is true if \(n=k\) i.e. you get to assume that \[(1+x)^k\geq 1+kx\]

OpenStudy (anonymous):

like after n=1 Assume true for some n=k after proving this so consider (1+x)^k + (1+x) ^k+1 is greater than or equal to 1+kx + (1+x)^k+1 where x>-1

OpenStudy (anonymous):

then using the fact that \[(1+x)^k\geq 1+kx\] prove that \[(1+x)^{k+1}\geq 1+(k+1)x\]

OpenStudy (anonymous):

it is not plus, it is times

OpenStudy (anonymous):

or was that the hint?

OpenStudy (anonymous):

really? cause for another problem my teacher showed original ...k^2 = (k(k+1) (2k+1))/6 then k^2 + (k+1) ^2 =( (k(k+1) (2k+1))/6 ) + (k+1)^2

OpenStudy (anonymous):

on that problem, i am assuming that the left hand side was a SUM right?

OpenStudy (anonymous):

\[\sum_{i=1}^k i^2=\frac{k(k+1)(2k+1)}{6}\]

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

that is way you got to add the last term you get to assume that \[\sum_{i=1}^k i^2=\frac{k(k+1)(2k+1)}{6}\] then note that \[\sum_{i=1}^{k+1} i^2=\sum_{i=1}^k i^2+(k+1)^2\]

OpenStudy (anonymous):

you have assumed that \[\sum_{i=1}^k i^2=\color{red}{\frac{k(k+1)(2k+1)}{6}}\] which makes \[\sum_{i=1}^{k+1} i^2=\sum_{i=1}^k i^2+(k+1)^2=\color{red}{\frac{k(k+1)(2k+1)}{6}}+(k+1)^2\]

OpenStudy (anonymous):

mhm

OpenStudy (anonymous):

ok in this case two things are different a) it is not a sum, it is a produce and b) it is not an equality, it is an inequality

OpenStudy (anonymous):

so we start by assuming that we know \[(1+x)^k\geq 1+kx\]

OpenStudy (anonymous):

since \(x>-1\) we know that \(1+x>0\) so the inequality stays the same if we multiply by \(1+x\) on both sides

OpenStudy (anonymous):

doing that, we get the inequality \[(1+x)^k(1+x)\geq (1+kx)(1+x)\] or \[(1+x)^{k+1}\geq (1+kx)(1+x)\]

OpenStudy (anonymous):

now the left side is just what we want lets see what we can do with the right side i.e. lets multiply out

OpenStudy (anonymous):

\[(1+x)^{k+1}\geq 1+kx+x+kx^2\] which we can rewrite as \[(1+x)^{k+1}\geq 1+(k+1)x+kx^2\]

OpenStudy (anonymous):

now we know that \(x^2\geq 0\) and so \(kx^2\geq 0\) making a string of inequalities \[(1+x)^{k+1}\geq 1+(k+1)x+kx^2\geq 1+(k+1)x\] ignoring the middle term, this is exactly what we want

OpenStudy (anonymous):

ohhhh

OpenStudy (anonymous):

not sure if any of that was clear, or if the goal was clear, but just to restate the goal, we are trying to show that if \[(1+x)^k\geq 1+kx\] then \[(1+x)^{k+1}\geq 1+(k+1)x\]

OpenStudy (anonymous):

tytytytyytyty i think i get it

OpenStudy (anonymous):

yw these are sometimes tricky to understand, but usually simple algebra proves them but you can't generalize from one proof to the next i.e. don't think they always work the same way summations are different than inequalities